OpenStudy (anonymous):

Factor x^4+21x^2-100

rishavraj (rishavraj):

clue 25 * -4 = 1-00 25 - 4 = 21

rishavraj (rishavraj):

i meant -100

OpenStudy (anonymous):

So I guess I could use the difference of squares?

rishavraj (rishavraj):

hmmmm i dont think so x^4 + 21x^2 - 100 = 0 x^4 - 4x^2 + 25x^2 - 100 = 0

OpenStudy (anonymous):

And then factor by grouping yes?

rishavraj (rishavraj):

yeah @PrincessDeliah